Determine the quadratic equation with integral coefficients having the roots: 3/2 and -1/2

Answer by ankor@dixie-net.com(22740)

Determine the quadratic equation with integral coefficients having the roots: 3/2 and -1/2
:
Obtain factors from the two roots
x = 3/2
2x = 3
(2x - 3) = 0
and
x = -1/2
2x = -1
(2x + 1) = 0
:
FOIL (2x-3)(2x+1) = 4x^2 + 2x - 6x - 3 = 4x^2 - 4x - 3